OpenStudy (anonymous):

Which of the following could correctly describe the graph of the equation below? Check all that apply. (x - 4)2 + (y + 3)2 = 36 A. Hyperbola opening left and right B. Circle C. Square D. Ellipse E. Parabola that opens up F. Hyperbola opening up and down

OpenStudy (anonymous):

General formula x^2 + y^2 = radius^2 is a circle So it is a circle however, a circle is a special kind of ellipse.. so I would say both a circle and an ellipse

OpenStudy (anonymous):

Are those 2's outside the parentheses indented to be exponents? If not, this is not any of those, it's a line. If so then it's a circle and an elipse

OpenStudy (anonymous):

If the equation of a parabola is y = 7(x - 4)2 + v, and its vertex is (4, -2), what is the value of v in the equation?

OpenStudy (radar):

It is a circle with center (-4,3) with a radius of 6

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Recall that the general form for a parabola is: \[y-y_0 = K(x-x_0)^2\] Where K is a multiplicative constant that changes the scale of the parabola, and \(x_0, y_0 \) are the coordinates of the vertex.

OpenStudy (anonymous):

in your case, moving the v to the otherside shows that \(y_0 = v\). So v would be the y value of your vertex.

OpenStudy (anonymous):

Which of the coordinate pairs below is a solution to the following system of equations? x2 + y2 = 157 x + y = 17

OpenStudy (anonymous):

Which of the coordinate pairs below is a solution to the following system of equations? x2 + y2 = 157 x + y = 17

OpenStudy (anonymous):

Its a real pleasure to see prof. Ram here